OpenStudy (anonymous):

determine the number of moles of squaric acid in 191.8 g

OpenStudy (anoop27):

191.8/114.06 =1.68

OpenStudy (kkutie7):

find its molecular formula first C4H2O4 use this to find the weight in gram per mole by using the masses in the periodic table \[4(C)+2(H)+4(O)= molar mass \] once you fin that you can do this \[191.8gC_{4}H_{2}O_{4}*\frac{1mole}{gmolarmass}\] this should give you the correct answer =)